OpenStudy (anonymous):

Solve 5x^2 = -30x - 65

OpenStudy (anonymous):

Bring the two terms on the right over to the left Then, factor out a 5 from each term. From there use factoring, or quadratic eq or completing the square to solve based in how your teacher wants you to solve. ^_^

OpenStudy (jmark):

Given Solve 5x^2 = -30x - 65 5x^2+30x+65=0 divide by 5 x^2+6x+13=0 use quadratic formula x=-3±2i ICSE class 9 question papers - http://icse.edurite.com/icse-question-papers/icse-question-papers-for-class-9.html